# EX-FEVER: A Dataset for Multi-hop Explainable Fact Verification

### v2

@@ Abstract
Fact verification aims to automatically probe the veracity of a claim based on several pieces of evidence. Existing works are always engaging in accuracy improvement, let alone explainability, a critical capability of fact verification systems. Constructing an explainable fact verification system in a complex multi-hop scenario is consistently impeded by the absence of a relevant, high-quality dataset. Previous datasets either suffer from excessive simplification or fail to incorporate essential considerations for explainability. To address this, we present EX-FEVER, a pioneering dataset for multi-hop explainable fact verification. With over 60,000 claims involving 2-hop and 3-hop reasoning, each is created by summarizing and modifying information from hyperlinked Wikipedia documents. Each instance is accompanied by a veracity label and an explanation that outlines the reasoning path supporting the veracity classification. Additionally, we demonstrate a novel baseline system on our EX-FEVER dataset, showcasing document retrieval, explanation generation, and claim verification, and validate the significance of our dataset. Furthermore, we highlight the potential of utilizing Large Language Models in the fact verification task. We hope our dataset could make a significant contribution by providing ample opportunities to explore the integration of natural language explanations in the domain of fact verification.
@@ 2.1 Fact Verification
Fact verification is a task similar to natural language inference, where the target is to predict whether evidence entails a claim. There are several models carefully designed for veracity reasoning based on retrieved evidence, including transformer-based methods Jiang et al. (2021); Kruengkrai et al. (2021) and graph-based methods Zhou et al. (2019); Liu et al. (2020); Zhong et al. (2020). Existing works can be roughly grouped into two categories aiming to enhance performance and robustness, respectively.
@@ 3.3 Overall Annotating Process
In order to harvest a high-quality dataset, we employed additional annotators as quality inspectors to scrutinize the annotations submitted by the primary annotators. Further details can be found in A.1.
@@ A.2 Details of the baseline system implementation
Verdict Predict We designate the claim to be verified as the MNLI task hypothesis and the relevant explanation of the verdict as the MNLI task premise. The BERT model is fine-tuned by using the Hugging Face Transformers library2. We implement the GEAR model Zhou et al. (2019), a graph-based text reasoning model, as a comparative alternative. The GEAR model conceptualizes sentences functioning either as evidence or claims within a graph theoretic framework, regarding such sentences as nodes on a graph. An evidential reasoning network and evidential aggregator transmit evidential information and make predictions. By combining these elements, GEAR leverages both evidence and claim in BERT to obtain an evidence representation $e_{i}$. The claim is fed into BERT alone to obtain representation $c$. Translating evidential and claim sentences into graph nodes expressing their relations, GEAR provides an integrative approach to adjudicating claim veracity via reasoning across the evidence.
@@
Check the claim: [claim] Choices:[’SUPPORT’, ’REFUTE’,’NOT ENOUGH INFO’] Answer:
Claim: [claim] Evidence: [golden documents] Evaluate the claim based on the provided evidence and choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’.
Claim: [claim] Evidence: [evidence] Evaluate the claim based on the provided evidence and choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’. Provide a brief explanation for your choice.
I will provide you with evidence and a claim. Your task is to determine if the claim is supported, refuted, or if there is not enough information based on the given evidence. You need to choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’. After choosing a label, please provide a brief explanation for your choice. Example 1: Evidence: [evidence] Claim: [claim] Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer: [answer] Explanation: [explanation] Example N: Evidence: [evidence] Claim: [claim] Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer: [answer] Explanation: [explanation] Now, please evaluate the following claim based on the provided evidence: Evidence: [evidence] Check the claim: + [claim] + from the above evidence Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer:

### v3

@@ Abstract
Fact verification aims to automatically probe the veracity of a claim based on several pieces of evidence. Existing works are always engaging in accuracy improvement, let alone explainability, a critical capability of fact verification systems. Constructing an explainable fact verification system in a complex multi-hop scenario is consistently impeded by the absence of a relevant, high-quality dataset. Previous datasets either suffer from excessive simplification or fail to incorporate essential considerations for explainability. To address this, we present EX-FEVER, a pioneering dataset for multi-hop explainable fact verification. With over 60,000 claims involving 2-hop and 3-hop reasoning, each is created by summarizing and modifying information from hyperlinked Wikipedia documents. Each instance is accompanied by a veracity label and an explanation that outlines the reasoning path supporting the veracity classification. Additionally, we demonstrate a novel baseline system on our EX-FEVER dataset, showcasing document retrieval, explanation generation, and claim verification, and validate the significance of our dataset. Furthermore, we highlight the potential of utilizing Large Language Models in the fact verification task. We hope our dataset could make a significant contribution by providing ample opportunities to explore the integration of natural language explanations in the domain of fact verification.  [^1] We make the EX-FEVER publicly accessible through https://github.com/dependentsign/EX-FEVER
@@ 2.1 Fact Verification
Fact verification is a task similar to natural language inference, where the target is to predict whether evidence entails a claim. There are several models carefully designed for veracity reasoning based on retrieved evidence, including transformer-based methods Jiang et al. (2021); Kruengkrai et al. (2021) and graph-based methods Zhou et al. (2019); Liu et al. (2020); Zhong et al. (2020); Xu et al. (2023, 2022); Gong et al. (2024). Existing works can be roughly grouped into two categories aiming to enhance performance and robustness, respectively.
@@ 3.3 Overall Annotating Process
To ensure the collection of a high-quality dataset, we engage additional annotators as quality inspectors to meticulously review the annotations submitted by the primary annotators. Each piece of collected data undergoes consistency checks by the annotators to guarantee quality. Further details can be found in A.1.
@@ 6 Conclusion
Acknowledgement
This work is supported by National Natural Science Foundation of China (62372454, 62141608)
@@ A.1 Details of data collection
Datasets Multi-hop Statement Count Explanation Type Class Count Source
HOVER 26,171 None 2 Wiki
FEVER 185,445 None 3 Wiki
e-FEVER - Machine Generated 3 Wiki
EX-FEVER 60,000 Human Annotated 3 Wiki
LIAR-PLUS - 12,836 Extracted Justification 6 Fact Check
PUBHEALTH - 11,832 Justification 4 Fact Check
Table 7: Comparison of Related Datasets. Note: e-FEVER dataset size is not revealed. LIAR-PLUS and PUBHEALTH are crawled from fact-check websites and are not practical for multi-hop classification
@@ A.2 Details of the baseline system implementation
Verdict Predict We designate the claim to be verified as the MNLI task hypothesis and the relevant explanation of the verdict as the MNLI task premise. The BERT model is fine-tuned by using the Hugging Face Transformers library3. We implement the GEAR model Zhou et al. (2019), a graph-based text reasoning model, as a comparative alternative. The GEAR model conceptualizes sentences functioning either as evidence or claims within a graph theoretic framework, regarding such sentences as nodes on a graph. An evidential reasoning network and evidential aggregator transmit evidential information and make predictions. By combining these elements, GEAR leverages both evidence and claim in BERT to obtain an evidence representation $e_{i}$. The claim is fed into BERT alone to obtain representation $c$. Translating evidential and claim sentences into graph nodes expressing their relations, GEAR provides an integrative approach to adjudicating claim veracity via reasoning across the evidence.
@@
Claim only Check the claim: [claim] Choices:[’SUPPORT’, ’REFUTE’,’NOT ENOUGH INFO’] Answer:
W/o explanation required Claim: [claim] Evidence: [golden documents] Evaluate the claim based on the provided evidence and choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’.
W/ explanation required Claim: [claim] Evidence: [evidence] Evaluate the claim based on the provided evidence and choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’. Provide a brief explanation for your choice.
Few-shot prompt I will provide you with evidence and a claim. Your task is to determine if the claim is supported, refuted, or if there is not enough information based on the given evidence. You need to choose one of the following labels: ’SUPPORT’, ’REFUTE’, or ’NOT ENOUGH INFO’. After choosing a label, please provide a brief explanation for your choice. Example 1: Evidence: [evidence] Claim: [claim] Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer: [answer] Explanation: [explanation] Example N: Evidence: [evidence] Claim: [claim] Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer: [answer] Explanation: [explanation] Now, please evaluate the following claim based on the provided evidence: Evidence: [evidence] Check the claim: + [claim] + from the above evidence Choices: [’SUPPORT’, ’REFUTE’, ’NOT ENOUGH INFO’] Answer: